2. Let X denote the number of times (1, 2, or 3 times) a certain machine malfunctions on any given day. Let Y denote the numbLet X denote the number of times (1, 2, or 3 times) a certain machine malfunctions on any given day.
Answer #1

(a) x takes values 1,2,3. Then, 0. (X ) = 0.05 +0.05 +0 - P(x - 2) = 0:05+0,1 +0.2=0-35 PLX = 3) 0.1t 0-35+ 0.13 0.55 P.M.F oMean of x = E() = 3 2 P(x= x) 221 = X 0+ 2 * 0.35 + 3X0.55 -.0.1+ 0.7+ 1.65 : 2.45 (6) y takes values 19213 Then, P(y=1) = 0.The s. Mor of 4 =P (4=4) 0.5 if y=2 0.3 if y=3Man of y = E(4) y Ply=y) = 10 2 + 2x0.5 + 3 0:3 © PY = 31= 2) PCY=3, *22) P(x=2) 0-2 0.35 = 0.571 a) E( XY) 3 22 3 4 7 (x = xX and Y no + are correlated We then know that if cov (x40 Here, COV( x4) = E(XY) - E( E(Y) = 5:15 - (2045) * 2.1 - 5.15 5.145